# Stochastic Block Covariance Matrix Estimation

### v1

@@ 4.2 Broader performance analysis
In our second experiment we pitted the block covariance estimation method against a variety of covariance estimators widely used in the literature, allowing for the true covariance matrix $\Sigma$ to deviate from the grouped covariance assumption. Again, we generated data according to the Gaussian model $\bm{y}_{i}\sim N(0,\Sigma)$ under various assumptions on $\Sigma$ commonly found in the literature. $\Sigma$ remained fixed across 100 replicates for each scenario. Figure 3 shows example $\Sigma$ matrices under each of these scenarios.
@@ 4.2.1 Experimental design
    MA(1): $~{}\sigma_{ij}=\rho^{|i-j|}\mathbf{1}\{|i-j|\leq 1\},$ $1\leq i,j\leq p$ (17)
    AR(1): $~{}\sigma_{ij}=\rho^{|i-j|},$ $1\leq i,j\leq p$ (18)
    Long-range: $~{}\sigma_{ij}=\frac{1}{2}[(|i-j|+1)^{2H}-2|i-j|^{2H}+(|i-j|-1)^{2H}],$ $1\leq i,j\leq p$ (19)
    Toeplitz: $~{}\sigma_{ij}=\rho|i-j|^{-(\alpha+1)},$ $1\leq i,j\leq p$ (20)
@@
Apart from these sparsity structures, we also considered lower-dimensional structures, following Ledoit and Wolf (2020), where covariance matrices were generated from eigen decomposition. Eigenvalues are either discrete (40%, 40%, and 20% of the eigenvalues were set to 10, 3, and 1, respectively), or uniformly distributed (eigenvalues were drawn uniformly between 1 and 10). Eigenvectors were randomly generated to form orthonormal matrices. Discrete eigenvalues tended to induce a block structure in the covariance matrices.
We also considered two other structures where the group covariance assumption is violated: (1) mixtures of block covariance matrices, where true covariance matrix was a mixture of three independently generated inhomogeneous grouped covariance matrices $(k^{*},\tau,\delta_{1},\delta_{2},\delta_{3})=(p/5,10,0.5,\allowbreak 0.2,0.3)$ with equal mixing weights. (2) random Inverse-Wishart distributed covariance matrices, where the true covariance matrix was drawn from $IW(\nu_{0},(\nu_{0}-p-1)\Sigma_{ig})$, centered at the previously described inhomogeneous grouped covariance matrix $\Sigma_{ig}$. The precision parameter $\nu_{0}$ controlled deviation from the group structure assumption, and we set $\nu_{0}=p+2$.
@@ 4.2.2 Competing estimators
We considered block covariance matrix estimators under three priors: weakly-informative, MLE-equivalent, and our proposed hierarchical prior. Each Bayesian method ran 5000 MCMC iterations, discarding the first 500 as burn-in, and every fifth posterior sample retained. The posterior mean served as the estimator for Bayesian methods.
@@
Banding estimator applies hard thresholds to off-diagonal entries of sample covariance matrix based on variables’ distance, setting distant covariances to zero. Tapering estimator gradually shrinks off-diagonal entries with increasing distance, resulting in a dense matrix with diminishing covariances. Hard thresholding and SCAD estimator apply thresholds to individual entries of the sample covariance matrix. Adaptive LASSO estimator penalizes entries unequally using data-dependent weights. Ledoit-Wolf linear shrinkage applies uniform shrinkage to all sample eigenvalues, pulling them toward a central value, while the nonlinear shrinkage estimator adjusts the shrinkage intensity based on the distribution of the sample eigenvalues, allowing for individualized shrinkage.
@@ 4.2.3 Comparison of estimators across various covariance structures
In Figure 4, the block covariance estimator using the hierarchical prior (lavender line) and weakly-informative prior (blue line) generally outperformed alternative estimators, except in cases involving order-dependent structures with rapidly decaying dependencies and randomly generated Inverse-Wishart covariances. This result highlighted the adaptability of the group covariance matrix assumption across diverse covariance types. Notably, the hierarchical prior consistently surpassed the weakly-informative prior, demonstrating its capacity to flexibly accommodate different structural patterns in the data.
For covariance with order-dependent sparsity, the banding/tapering estimator consistently performed better than other estimators, except in cases with long-range dependencies. This is because the banding/tapering estimator correctly bet on the underlying structure, whereas other estimators struggled due to misspecification. Moving from left to right across the first row of panels, as dependency decayed more slowly, covariance matrices had less sparsity and larger off-diagonal values (as shown in Figure 3). In such cases, the assumptions underlying the banding/tapering and thresholding estimators became less appropriate, the block covariance estimator performed better, particularly as long-range dependencies grew stronger.
@@
Across all scenarios, our estimator outperformed the sample covariance matrix. Notably, the weakly-informative prior sometimes underperformed relative to the sample covariance matrix, particularly when $n>p$ and for Inverse-Wishart or rapidly decaying, order-dependent structures. This result illustrated the limitations of a fixed shrinkage target and the necessity of adjustable shrinkage intensity. The hierarchical prior provided adaptability in shrinkage, ensuring robust performance even under challenging conditions.
@@ 5 Relation to Other Covariance Estimation Methods
Stochastic block covariance estimation is inherently a regularized factor model, enforcing identical factor loadings within the same cluster. It also inherits the rationale behind Stein’s class of shrinkage estimators, which reduces the dispersion of sample eigenvalues through clustering. Discussing clustered factor model and Stein’s estimators can mirror stochastic block covariance estimation from different perspectives, enhancing our understanding on how it achieves strong performance across various covariance structures.
@@ 5.1 Clustered factor model
The block covariance matrix has several advantages over the clustered factor model. First, block covariance models are more general, as the clustered factor model only represents a subset of block covariance matrices that are positive definite (see Theorem 1 and Lemma 1 in Tong and Hansen (2023)). Second, clustered factor models require $k\times r+r$ parameters, where $r$ is the number of factors, leading to higher parameter complexity unless the block covariance matrix is highly parsimonious ($k<2r-3$). Third, clustered factor model requires selecting an appropriate number of factors, and the number of factors is bound by the covariance structure, posing a great challenge in model estimation. One might consider an infinite factor model and introduce a prior to truncate it to a finite number (Bhattacharya and Dunson, 2011). However, undoubtedly, estimating the number of factors would introduce additional complexity and computational burden for block covariance estimation. Additionally, block covariance models are more straightforward and transparent for interpretation, as latent factors and their loadings can often be challenging to interpret meaningfully.
@@ 5.2 Stein’s class of shrinkage estimators
Upon revisiting the numerical experiment results shown in Figures 4 and Figure 14, the consistently good performance of the Ledoit-Wolf shrinkage estimator captured our attention. The Ledoit-Wolf shrinkage estimator consistently ranked as the second-best method, following closely behind the block covariance matrix estimator in most cases, except for those involving banded/tapered or sparse banded structures. It outperformed or remained competitive with the block covariance estimators when the true covariance matrix was generated from an inverse-Wishart distribution or followed an MA(1) or AR(1) process in settings where $n>p$. Notably, it consistently surpassed the sample covariance matrix even in challenging cases.
@@
A key development in extending Stein’s estimator to high dimensional settings stems from advances in random matrix theory, particularly the Marčenko-Pastur equation (Wigner, 1955; Marčenko and Pastur, 1967; Silverstein, 1995). Building upon these fundamental results, various nonlinear shrinkage estimators have been developed, with the Ledoit-Wolf shrinkage estimators (Ledoit and Wolf, 2022) being notable examples.
The block covariance estimator shares some conceptual similarities with Stein’s estimators, since both reduce the dispersion of sample eigenvalues, albeit through different mechanisms. While Stein’s approach shrinks eigenvalues directly, guided by the rotation of the sample eigenvectors, the block covariance estimator achieves shrinkage by clustering eigenvalues, driven by a block-sparse rotation, effectively eliminating within-group variance.
@@
Figure 5 displays a comparison of estimation accuracy between block covariance estimator (BCM) and Stein’s class of estimators across various covariance structures.  [^1] To prevent numerical instability in density estimation, we truncated sample eigenvalues below $10^{-8}$ to zero when applying the Ledoit-Wolf method. The block covariance estimators generally outperformed or was comparable to the oracle estimators, except when the covariance matrix was generated from an MA(1) process (characterized by rapid decay in dependency) or a random inverse-Wishart distribution. Across nearly all scenarios, the block covariance estimator consistently surpassed Ledoit-Wolf estimator. This advantage arose from the BCM’s reliance on rotation learning, which effectively leveraged structural information within the data, allowing better adaptation when such structure was present. In contrast, Stein’s family estimators operated without strong structural assumptions, making them overly conservative and limiting their flexibility and adaptability to various covariance matrix configurations.
The block covariance estimator and Stein’s estimator originate from two classical models: factor analysis and principle component analysis (PCA), respectively. Our model introduces additional ergularization on factor loadings by enforcing identical factor loadings within the same cluster while also promoting sparsity. Similar extensions of factor analysis and PCA can be found in Rohe and Zeng (2023), where varimax rotation is recommended to promote sparsity and interpretability for non-Gaussian data.
@@ 6.1 Neural coordination
Previous studies (Caruso et al., 2018; Jun et al., 2022; Schmehl et al., 2024; Chen et al., 2024) have shown that individual neurons can encode two simultaneously presented stimuli through a turn-taking dynamics called "code juggling", where neurons randomly switch between two composite stimuli either at a slow rate across trials or at a fast rate within a trial. However, this alternation between stimuli is not entirely random. Some evidence suggests a coordinated pattern of code juggling across neural populations (Jun et al., 2022; Schmehl et al., 2024). This coordination can be represented by a block structure in covariance matrices, which may provide valuable insights into how neural circuits contribute to perception and behavior, a fundamental question in neuroscience.
Figure 6: Score matrix of 74 neurons across 23 trials, with each entry showing the neuron’s tendency to encode stimulus A (blue) or B (red) during simultaneous dual-stimulus exposure. Co-fluctuating and oppositely fluctuating patterns among neurons suggest clustering in the covariance structure.
@@
We analyze a dataset from Ruff and Cohen (2016), where monkeys performed a motion direction change detection task while neural activity was recorded. In trials involving dual stimuli (A and B), adjacent drifting Gabor patches with different orientations were presented. In single stimulus trials, only one of the constituent drifting Gabor patches was visible. A multi-electrode array is planted in V1 area recording spike activities of 101 neurons simultaneously. We use a subset of the data and preprocess it to generate a score matrix for 74 neurons across 23 trials, where each entry reflects a neuron’s inclination to encode stimulus A. For a detailed description of the preprocessing procedure, refer to Appendix F.1.
Figure 6 displays the score matrix, organized according to clustering from our block covariance matrix estimator. It reveals distinct clusters of neurons with different patterns in their activity fluctuations. Our primary focus is on examining the block structure of the covariance matrix underlying these fluctuations. Figure 7 and 8 show the posterior mean of covariance matrices and posterior similarity matrices, which represent the posterior probability of two neurons being clustered together. Neurons are sorted based on the clustering results from the weakly informative prior to facilitate comparison.
Figure 7: Posterior mean of covariance matrices under different priors in the neural population coordination application. The weakly informative prior oversimplifies the block structure and misrepresents between-block covariances, while the MLE-equivalent prior favors sparsity, resulting in many isolated small-sized blocks. The hierarchical prior, with adaptive shrinkage, refines the block structure to better reflect the underlying data.
@@
In Figure 7, three primary clusters emerge: a median-sized cluster of 15 neurons and two larger clusters of around 26 and 33 neurons. Neurons within the first cluster are highly interconnected, positively correlated with the second cluster and negatively correlated with the third.
The weakly informative prior excessively shrinks the covariance toward a block diagonal structure, oversimplifying the block structure and misrepresenting between-block covariances. As shown in Figure 7, the weakly informative prior manifests two main issues in covariance estimation: (1) diminished within-block covariance, where some high positive within-block covariances among certain neurons are significantly reduced due to the shrinkage. (2) a reversal of the sign of between-block covariance, where some negative covariances between the second and third clusters are shrunk to weak positive covariances. These issues are reflected in its underlying block assignments, which are primarily influenced by within-block covariances and fail to account for the heterogeneity in between-block covariances. For instance, Figure 7 reveals a subgroup of neurons in the second block that are positively correlated with the first block and negatively correlated with the third block, while other neurons in the second block display the opposite pattern. However, this nuanced sub-block structure is not clearly captured by the weakly informative prior, as seen in the posterior similarity matrix in Figure 8.
@@
The MLE-equivalent prior focuses on capturing the sparsity of the sample covariance matrix, resulting in the formation of too many blocks, particularly isolated small-sized ones. In Figure 7, the numerous white ribbons in the covariance estimator indicate that many entries are overly shrunk toward zero, especially within small blocks. Comparing the posterior similarity matrix in Figure 8 with the sample covariance matrix in Figure 7, it appears that the MLE-equivalent prior prioritizes sparsity over preserving the block structure. The block assignment is largely driven by its sparsity, rather than by meaningful block patterns. Additionally, the posterior similarity matrix of the MLE-equivalent prior shows a concentrated distribution, which may suggest slow MCMC mixing or a tendency to get stuck in a local mode.
@@
Both the weakly informative and MLE-equivalent priors suffer from the mis-specification of the shrinkage target and lack adaptive intensity. In contrast, the hierarchical prior refines the block structure to better align with the data through adaptive shrinkage. As seen in Figure 7, the hierarchical prior is sensitive to the heterogeneity of between-block covariances, capturing a more detailed and accurate structure than the weakly informative prior. For instance, it effectively captures sub-block structures within the second and third large blocks. Unlike the MLE-equivalent prior, the hierarchical prior avoids introducing isolated small blocks and excessive shrinkage toward zero, providing a smoother, more accurate estimator. Figure 8 further illustrates how the hierarchical prior enhances cluster segmentation and restores essential features. This refined block structure affirms the presence of coordination within the neural population, providing a clear depiction of potential neural circuits.

### v2

@@ 4.2.1 Experimental design
In our second experiment we pitted the block covariance estimation method against a variety of covariance estimators widely used in the literature, allowing for the true covariance matrix $\Sigma$ to deviate from the grouped covariance assumption. Again, we generated data according to the Gaussian model $\bm{y}_{i}\sim N(0,\Sigma)$ under various assumptions on $\Sigma$ commonly found in the literature. The true $\Sigma$ remained fixed across 100 replicates for each scenario, which are shown in Figure 3 corresponding to the sparsity assumptions listed below:
@@
    MA(1): $~{}\sigma_{ij}=\rho^{|i-j|}\mathbf{1}\{|i-j|\leq 1\},$ (17)
    AR(1): $~{}\sigma_{ij}=\rho^{|i-j|},$ (18)
    Long-range: $~{}\sigma_{ij}=\frac{1}{2}[(|i-j|+1)^{2H}-2|i-j|^{2H}+(|i-j|-1)^{2H}],$ (19)
    Toeplitz: $~{}\sigma_{ij}=\rho|i-j|^{-(\alpha+1)}.$ (20)
@@
Apart from these sparsity structures, we also considered lower-dimensional structures, following Ledoit and Wolf (2020), where covariance matrices were generated from eigen decomposition. Eigenvalues were either taken to be discrete (40%, 40%, and 20% of the eigenvalues were set to 10, 3, and 1, respectively), or drawn uniformly distributed between 1 and 10. Eigenvectors were randomly generated to form orthonormal matrices. Discrete eigenvalues tended to induce a block structure in the covariance matrices.
We also considered two other structures where the group covariance assumption was violated: (1) mixtures of block covariance matrices, where true covariance matrix was a mixture of three independently generated inhomogeneous grouped covariance matrices $(k^{*},\tau,\delta_{1},\delta_{2},\delta_{3})=(p/5,10,0.5,\allowbreak 0.2,0.3)$ with equal mixing weights. (2) random Inverse-Wishart distributed covariance matrices, where the true covariance matrix was drawn from $IW(\nu_{0},(\nu_{0}-p-1)\Sigma_{ig})$, centered at the previously described inhomogeneous grouped covariance matrix $\Sigma_{ig}$. The precision parameter $\nu_{0}$ controlled deviation from the group structure assumption, and we set $\nu_{0}=p+2$.
@@ 4.2.2 Competing estimators
We considered block covariance matrix estimators under three priors: weakly-informative, MLE-equivalent, and our proposed hierarchical prior. For each method, estimates were obtained with 5000 Markov chain Monte Carlo iterations, discarding the first 500 as burn-in, and every fifth posterior sample retained. The posterior mean served as the estimator for Bayesian methods.
@@
The banding estimator applies hard thresholding to off-diagonal entries of the sample covariance matrix based on the distance between the positions of the variables within a user-specified ordering, setting distant covariances to zero. These are most appropriate for time series data where a linear ordering is known and is taken to provide information about relative strength of correlation between variables. The tapering estimator is similar to the banding estimator but instead of hard thresholding, it gradually shrinks off-diagonal entries with increasing distance, resulting in a dense matrix with diminishing covariances. Hard thresholding and SCAD estimators apply thresholds to individual entries of the sample covariance matrix – no particular order of presentation is assumed. The adaptive LASSO estimator penalizes entries unequally using data-dependent weights. Ledoit-Wolf linear shrinkage applies uniform shrinkage to all sample eigenvalues, pulling them toward a central value, while the nonlinear shrinkage estimator adjusts the shrinkage intensity based on the distribution of the sample eigenvalues, allowing for individualized shrinkage.
@@ 4.2.3 Comparison of estimators across various covariance structures
For covariance with order-dependent sparsity, the banding/tapering estimator consistently performed better than other estimators, except in cases with long-range dependencies. This is because the banding/tapering estimators correctly bet on the underlying structure, whereas other estimators struggled due to misspecification. Moving from left to right across the first row of panels, as dependency decayed more slowly, covariance matrices had less sparsity and larger off-diagonal values (as shown in Figure 3). In such cases, the assumptions underlying the banding/tapering and thresholding estimators became less appropriate, the block covariance estimator performed better, particularly as long-range dependencies grew stronger.
@@
Across all scenarios, our estimator outperformed the sample covariance matrix. Notably, the weakly-informative prior sometimes underperformed relative to the sample covariance matrix, particularly when $n>p$ and for Inverse-Wishart or rapidly decaying, order-dependent structures. This result illustrates the limitations of a fixed shrinkage target and the necessity of adjustable shrinkage intensity. The hierarchical prior appears to provide adaptability in shrinkage, ensuring robust performance even under challenging conditions.
@@ 5 Relation to Other Covariance Estimation Methods
At this point, a comprehensive mathematical explanation remains elusive as to why the grouped covariance assumption works so well across various challenging scenarios. Here we shed further light on the the underlying theory by drawing attention to other theoretical frameworks commonly adopted in the literature which seem to have close relations to the stochastic block covariance model.
@@ 5.1 Clustered factor model
In spite of the similarity to stochastic block covariance model, as an estimation framework, the clustered factor model is more restrictive and less efficient in its representation. First, it represents a subset of stochastic block covariance matrices that are positive definite (see Theorem 1 and Lemma 1 in Tong and Hansen (2023)), and usually requires more parameters ($k\times r+r$) than stochastic block covariance matrices, except in cases where the latter is highly parsimonious ($k<2r-3$). Additional challenges present in deciding on an optimal number of factors and in interpreting the estimated factors and their loadings. Stochastic block covariance representation, in contrast, provides a more straightforward estimation and transparent interpretation.
@@ 5.2 Stein’s class of shrinkage estimators
In our experiments reported above, the Ledoit-Wolf shrinkage estimator consistently ranked as the second-best method, following closely behind the block covariance matrix estimator in most cases, except for those involving banded/tapered or sparse banded structures. It outperformed or remained competitive with the block covariance estimators when the true covariance matrix was generated from an inverse-Wishart distribution or followed an MA(1) or AR(1) process in settings where $n>p$. Notably, it consistently surpassed the sample covariance matrix even in challenging cases. We offer here a preliminary analysis of the similarities and differences between the Ledoit-Wolf estimator and our stochastic block covariance estimator.
@@
We note that the block covariance estimator shares some conceptual similarities with Stein’s estimators, since both reduce the dispersion of sample eigenvalues, albeit through different mechanisms. While Stein’s approach shrinks eigenvalues directly, guided by the rotation of the sample eigenvectors, the block covariance estimator achieves shrinkage by clustering eigenvalues, driven by a block-sparse rotation, effectively eliminating within-group variance.
@@
Figure 5 displays a comparison of estimation accuracy between stochastic block covariance estimator (BCM) and Stein’s class of estimators across various covariance structures.  [^1] To prevent numerical instability in density estimation, we truncated sample eigenvalues below $10^{-8}$ to zero when applying the Ledoit-Wolf method. The block covariance estimators generally outperformed or was comparable to the oracle estimators, except when the covariance matrix was generated from an MA(1) process (characterized by rapid decay in dependency) or a random inverse-Wishart distribution. Across nearly all scenarios, the stochastic block covariance estimator consistently surpassed Ledoit-Wolf estimator. This advantage arose from the BCM’s reliance on rotation learning, which effectively leveraged structural information within the data, allowing better adaptation when such structure was present. In contrast, Stein’s family estimators operated without strong structural assumptions, making them overly conservative and limiting their flexibility and adaptability to various covariance matrix configurations.
The stochastic block covariance estimator and Stein’s estimator originate from two classical models: factor analysis and principle component analysis (PCA), respectively. Our model introduces additional regularization on factor loadings by enforcing identical factor loadings within the same cluster while also promoting sparsity. Similar extensions of factor analysis and PCA can be found in Rohe and Zeng (2023), where varimax rotation is recommended to promote sparsity and interpretability for non-Gaussian data.
@@ 6.1 Neural coordination
Previous studies (Caruso et al., 2018; Jun et al., 2022; Schmehl et al., 2024; Chen et al., 2024) have shown that individual neurons can encode two simultaneously presented stimuli through a turn-taking dynamics called "code juggling", where neurons stochastically switch responding to two composite stimuli, such as at either a slow rate across trials or a fast rate within a trial. However, this alternation between stimuli is not entirely random. Some evidence suggests a coordinated pattern of code juggling across neural populations (Jun et al., 2022; Schmehl et al., 2024). This coordination can be represented by a block structure in covariance matrices, which may provide valuable insights into how neural circuits contribute to perception and behavior, a fundamental question in neuroscience.
Figure 6: (a) Score matrix of 74 neurons across 23 trials, with each entry showing the neuron’s tendency to encode stimulus A (purple) or B (red) during simultaneous dual-stimulus exposure. Co-fluctuating and oppositely fluctuating patterns among neurons suggested clustering in the covariance structure. (b) Posterior mean of covariance matrices under different priors. Neurons in (a) and (b) were sorted in the same order based on clustering results from stochastic block covariance estimation under the weakly-informative prior and the hierarchical prior. Black boxes outline the three main groups (size 15, 24, and 35) clustered under the weakly-informative prior. Within these black boxes, orange boxes represent sub-group structure detected by the hierarchical prior but ignored by the weakly-informative prior. The weakly informative prior oversimplified the block structure (orange) and misrepresented between-block covariances (purple), while the MLE-equivalent prior favored sparsity, resulting in many isolated small-sized blocks shown as white ribbons. The hierarchical prior, with adaptive shrinkage, refined the block structure to better reflect the underlying data.
@@
We analyzed a dataset from Ruff and Cohen (2016), where monkeys performed a motion direction change detection task while neural activities were recorded. In trials involving dual stimuli (both A and B), adjacent drifting Gabor patches with different orientations were presented. In single stimulus trials, only one of the constituent drifting Gabor patches was visible. A multi-electrode array was planted in V1 area recording spike activities of 101 neurons simultaneously. We used a subset of the data and preprocessed it to generate a score matrix for 74 neurons across 23 trials, where each entry reflects a neuron’s inclination to encode stimulus A. For a detailed description of the preprocessing procedure, refer to the Appendix F.1.
Figure 6(a) displays the score matrix, revealing distinct clusters of neurons with different patterns in their activity fluctuations. Examining the underlying block structure of these fluctuations was our primary focus. Figure 6(b) shows posterior mean of covariance matrices comparing shrinkage effects under different choices of hyperparameters. Figure 14 in the Appendix presents the posterior similarity matrix, which quantifies the posterior probability of each neuron pair being grouped together. Neurons in these three figures were arranged in the same order. Black boxes outline the three main groups identified under the weakly-informative prior. Within these groups, orange boxes highlight detailed sub-groups captured by the hierarchical prior but overlooked by the weakly-informative prior. Purple boxes highlight covariances that were misrepresented by the weakly-informative prior but well captured by the hierarchical prior.
To facilitate interpretation, special care is required for neuron reordering. For better comparison among different choices of hyperparameters, we incorporated clustering information from both the weakly-informative prior and the hierarchical prior using a two-stage procedure. We first reordered neurons based on hierarchical clustering, where the distance matrix was defined as 1 minus the mean of posterior covariance matrices under the weakly-informative prior. In this process, three major blocks emerged, outlined by black boxes in Figure 6(b). We then repeated the same procedure to reorder neurons within each block, but with respect to mean of posterior covariance matrices under the hierarchical prior instead of the weakly-informative prior. A finer structure ignored by the weakly-informative prior was revealed by the hierarchical prior, highlighted by orange and purple boxes in Figure 6(b).
@@
In Figure 6(b), three primary clusters emerged: a median-sized cluster of 15 neurons and two larger clusters of around 24 and 35 neurons, as outlined by black boxes. Neurons within the first cluster were highly correlated, positively correlated with the second cluster and negatively correlated with the third.
The weakly informative prior excessively shrank the covariance toward a block diagonal structure, oversimplifying the block structure and misrepresenting between-block covariances (particularly negative ones). As shown in Figure 6(b), the weakly informative prior manifested two main issues in covariance estimation: (1) diminished within-block covariances, where some high positive within-block covariances among certain neurons were significantly reduced due to the shrinkage (see orange boxes). (2) a reversal of the sign of between-block covariances, where some negative covariances were shrunk to weak positive covariances (see purple boxes). The large bias in covariance estimation was related to the oversimplified block structure estimation (see Figure 14). The weakly informative prior underestimate of the block count $k$, weakening the effective use of structural information on covariance estimation.
@@
The MLE-equivalent prior focused on capturing the sparsity of the sample covariance matrix, resulting in the formation of too many blocks, particularly isolated small-sized ones. In Figure 6(b), the numerous white ribbons in the covariance estimator indicated that many entries were overly shrunk toward zero, especially within small blocks. Comparing the posterior similarity matrix in Figure 14 with the sample covariance matrix in Figure 6(b), it appeared that the MLE-equivalent prior prioritized sparsity over preserving the block structure. The block assignment was largely driven by its sparsity, rather than by meaningful block patterns. Additionally, the posterior similarity matrix of the MLE-equivalent prior showed a concentrated distribution, which may suggest slow MCMC mixing or a tendency to get stuck in a local mode.
@@
Both the weakly informative and MLE-equivalent priors suffered from the mis-specification of the shrinkage target and lack adaptive intensity. In contrast, the hierarchical prior refined the block structure to better align with the data through adaptive shrinkage (see Figure 6). Unlike the weakly informative prior, the hierarchical prior was sensitive to the heterogeneity of between-block covariances (purple boxes), and effectively captures finer sub-block structures (orange boxes). Unlike the MLE-equivalent prior, the hierarchical prior avoided introducing isolated small blocks and excessive shrinkage toward zero, providing a smoother, more accurate estimator. Figure 14 further illustrated how the hierarchical prior enhanced cluster segmentation and restored essential features. This refined block structure affirmed the presence of coordination within the neural population, providing a clear depiction of potential neural circuits.